OpenStudy (anonymous):

What is the common ratio of the following sequence? 7, 21, 63, 189

mathslover (mathslover):

If i am right then it is 21 / 7 = 3 ..... 63 / 21 = 3

mathslover (mathslover):

Hence the common ratio of the following sequence will be 3

OpenStudy (lgbasallote):

\[r = \frac{a_2}{a_1}\] that just means common ratio is second term divided by first term

mathslover (mathslover):

It is called geometric progression .. got it @alexzndrac ?????

OpenStudy (anonymous):

okay thanks you guys

mathslover (mathslover):

No problem best of luck :)

mathslover (mathslover):

\[\large{\textbf{Welcome to openstudy}}\] .. http://openstudy.com/code-of-conduct for more information .. thanks!!!

OpenStudy (anonymous):

\[7 \times 3^{x}\] is the equation btw.